# Gaudin models and moduli space of flower curves

### v1

@@ 5.5. Compactification of the family of trigonometric Gaudin subalgebras
[…]
Figure 3. This is a curve $C\in\overline{M}_{6+2}$. We have $n=6$, $m=4$, and =($\mathcal{B}=(\{0\},\{1,2\},\{3\},\{4,5,6\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}^{trig}_{\theta}(C)=\Delta^{\mathcal{B}^{\prime}}(\mathcal{A}^{trig}_{\theta}(w_{1},w_{2},w_{3}))\mathcal{A}(z_{1},z_{2})\mathcal{A}(z_{4},z_{6},z_{6})$.
@@
The statement about Hilbert-Poincaré series follows by multiplying the Hilbert-Poincaré series of the tensor factors and dividing by the Hilbert-Poincaré series of $Z^{\prime}$. ∎
@@
Figure 4. This is a curve $C\in\overline{M}_{6+2}$ where $z_{0}$ does not lie on the same component as $z_{6+1}$. We have $n=6$, $m=4$, and =($\mathcal{B}=(\{0,1,2,3\},\{4\},\{5\},\{6\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}^{trig}_{\theta}(C)=\psi_{\theta}(j_{\mathcal{B}}^{6+1}\mathcal{A}(0,z_{1},z_{2},z_{3})))\psi_{\theta}(\Delta^{\mathcal{B}}(\mathcal{A}(0,z_{4},z_{5},z_{6}))$.
@@
We use the same idea as in the proof of Theorem 5.5 to show that the above coefficients of (14) and (15) are algebraically independent, and this would be sufficient. Namely, we consider the filtration on $S(\mathfrak{g})^{\otimes n}$ where the $n$-th component has degree $1$ and all the others have degree 00, take the leading terms with respect to this filtration, and show that the differentials of these leading terms at the point $y=(h,\ldots,h,\underbrace{f,\ldots,f}_{m-1},f+h)$ are linearly independent.
@@ 6.2. Compactification of the family of inhomogeneous Gaudin subalgebras
We will now study the compactification of the family of inhomogeneous Gaudin subalgebras with fixed $\chi\in\mathfrak{g}^{reg}$. This family will be parametrized by the space $\overline{F}_{n}$ of cactus flower curves from section 3.2. Recall that $\overline{F}_{n}$ is a compatification of $F_{n}=\mathbb{C}^{n}\setminus\Delta/\mathbb{C}$. For $(z_{1},\dots,z_{n})\in F_{n}$, we have the inhomogeneous Gaudin subalgebra $\mathcal{A}_{\chi}(z_{1},\dots,z_{n})$.
@@
Figure 5. This is a curve $C\in\overline{F}_{8}$. We have $n=8$, $m=3$, and ${\mathcal{S}}=(\{1,2,3,4\},\{5\},\{6,7,8\})$. The curve $C_{1}\in F_{4}$ has =($\mathcal{B}=(\{1\},\{2,3,4\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}_{\chi}(C)=\Delta^{\mathcal{B}}(\mathcal{A}_{\chi}(w_{1},w_{2}))\mathcal{A}(z_{2},z_{3},z_{4})\otimes\mathcal{A}_{\chi}\otimes\mathcal{A}_{\chi}(z_{6},z_{7},z_{8})$.
@@ 6.3. Conjectural closure for arbitrary $\chi\in\mathfrak{h}$
$F_{n,\mathfrak{g}}:=(\mathbb{C}^{n}\setminus\Delta/\mathbb{C})\times\mathfrak{h}^{reg}/\mathbb{C}^{\times}$
@@
For $n=1$ this Conjecture is [HKRW, Theorem 10.8]. Also, the particular case of this Conjecture with $n=2$ and $\mathfrak{g}_{1}$ being a root $\operatorname{\mathfrak{sl}}_{2}$ is discussed in [HKRW, Proposition 10.17]. Moreover, we expect that for $\mathfrak{g}=\operatorname{\mathfrak{sl}}_{m}$ all such spaces $\overline{F}_{n,\mathfrak{g}}$ are given by the construction of Bottman and Oblomkov [BO] of a moduli space of marked vertical lines in $\mathbb{C}^{2}$ with $n-1$ vertical lines without marked points and another vertical line with $m$ marked points [^4] We thank Alexei Oblomkov for pointing out this conjectural relation..
@@ 7.1. Degeneration of trigonometric to inhomogeneous
Recall ${\mathcal{F}}_{n}:={\mathbb{G}}^{n}\setminus\Delta/{\mathbb{G}}$, as defined in section 3.3. Our goal in this section is to define a family of subalgebras of $U\mathfrak{g}^{\otimes n}$ parametrized by $\mathcal{F}_{n}$ and to prove the following result.
@@
    After identifying $U_{\hbar}\mathfrak{g}(0)=S(\mathfrak{g})$ and $U_{\varepsilon}\mathfrak{g}=U\mathfrak{g}$, for $\varepsilon\neq 0$, this map equals $\operatorname{ev}_{\infty,z_{1},\dots,z_{n}}$ if $\varepsilon=0$ and $\operatorname{ev}_{\varepsilon^{-1},z_{1},\dots,z_{n}}$ otherwise.
@@
[…]
@@ 7.2. Quadratic parts of inhomogeneous/trigonometric Gaudin subalgebras
${\mathcal{F}}_{n}\rightarrow{\mathrm{Gr}}_{\mathbb{A}^{1}}(2n+1,U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n})\quad({\underline{z}},\varepsilon)\mapsto A^{\varepsilon}({\underline{z}})^{(2)}$
${\mathcal{F}}_{n}\rightarrow{\mathrm{Gr}}_{\mathbb{A}^{1}}(2n,\mathbb{C}[\hbar]\otimes U\mathfrak{g}^{\otimes n})\quad({\underline{z}},\varepsilon)\mapsto A^{\varepsilon}_{\chi}({\underline{z}})^{(2)}$
@@
¯$\overline{\mathcal{F}}_{n}\rightarrow{\mathrm{Gr}}_{\mathbb{A}^{1}}(2n+1,U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n})\quad\overline{\mathcal{F}}_{n}\rightarrow{\mathrm{Gr}}_{\mathbb{A}^{1}}(2n,\mathbb{C}[\hbar]\otimes U\mathfrak{g}^{\otimes n})$
@@
In section 7.1, we constructed $\mathbfcal A$, a family of subalgebras of $U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$ parametrized by ${\mathcal{F}}_{n}$, and $\mathbfcal A_{\chi}$, a family of subalgebras of $\mathbb{C}[\hbar]\otimes U\mathfrak{g}^{\otimes n}$ also parametrized by ${\mathcal{F}}_{n}$ (in both cases regarded over the base ring $\mathbb{C}[\hbar]$). Recall also that in [IKLPR], we constructed a (relative) compactification ¯$\overline{\mathcal{F}}_{n}$, see section 3.3.
@@
These extend to flat families of subalgebras faithfully parametrized by ¯$\overline{\mathcal{F}}_{n}$. Thus ¯$\overline{\mathcal{F}}_{n}$ is the compactified parameter space for these inhomogeneous/trigonometric subalgebras.
As in the proof of Theorem 5.17, it suffices to show that the family extends to a flat family parametrized by ¯$\overline{\mathcal{F}}_{n}$. Subsequently, the faithfulness of the parametrization follows from Corollary 7.10.
@@
    Each ${\mathcal{W}}_{\tau}$ is a union of strata of ¯$\overline{\mathcal{F}}_{n}$. These strata are constructed in an “operadic manner” from smaller ¯$\overline{\mathcal{F}}_{r}$ and $\overline{M}_{k+1}$. We prove that the restrictions of our generators to these strata are compatible with the generators of the Gaudin subalgebras over these strata.
@@
On the other hand, the sum of residues of $f(w)$ at all $w_{i}$ and $\infty$ is 00, so we have
@@
Let $\tau$ be a binary rooted forest. Let $v$ be an internal vertex of $\tau$. Let $p$ be the rightmost leaf of the left branch growing from $v$ and let $q$ be the leftmost leaf on the right branch growing from $v$. We make an affine change of the variable $u$ taking $z_{p}$ to 00 and $z_{q}$ to $1$; the new variable is
@@
Now, suppose that $v$ is a root of $\tau$. Let $p$ be the rightmost leaf on the tree growing from $v$. We make an affine change of the variable $u$ taking $z_{p}$ to 00 and preserving $\hbar^{-1}$; the new variable is
@@
In [IKLPR, §6.3], we considered a stratification of ¯$\overline{\mathcal{F}}_{n}$ indexed by pairs of set partitions; on each stratum certain of the $\nu_{ij}$ coordinates are equal to 0 and others are equal to $\infty$ (as $\delta_{ij}=\nu_{ij}^{-1}$, this is equivalent to setting some $\delta_{ij}=0$).
Let $A$ be a codimension 1 stratum of ¯$\overline{\mathcal{F}}_{n}$ such that $A\cap{\mathcal{W}}_{\tau}\neq\emptyset$. From [IKLPR, Remark 6.13], we have the following 2 possibilities:
@@
    $A=\overline{\mathcal{F}}_{I}\times_{\mathbb{A}^{1}}\overline{\mathcal{F}}_{J}$, determined by a partition $[n]=I\sqcup J$ compatible the forest $\tau$, by putting $\nu_{ji}=0$ for all $i\in I,\ j\in J$;
@@
$s_{l,v,\tau}^{m}=\begin{cases}j_{\{0\}\cup I}^{n+1}(s_{l,v,\tau_{I}}^{m})\quad\text{ if $v$ is in the subforest $\tau_{I}$}\\ \Delta^{\mathcal{B}}(s_{l,v,\tau_{J}}^{m})\quad\text{ otherwise.}\end{cases}$
@@
Now it remains to show that the above generators stay algebraically independent. This follows from the particular case of Knop’s theorem [Kn] (see [R3, Lemma 3.10] for more details), namely $[U\mathfrak{g}^{\otimes I}]^{\mathfrak{g}}$ is a free $\Delta(ZU\mathfrak{g})$-module and that the homomorphism $\Delta\cdot\operatorname{id}:U\mathfrak{g}\otimes[U\mathfrak{g}^{\otimes I}]^{\mathfrak{g}}\to U\mathfrak{g}^{\otimes I}$ factors as
@@
[…]
@@
$s_{l,v,\tau}^{m}=\begin{cases}j_{\{0\}\cup I}^{n+1}(s_{l,v,\tau_{I}}^{m})\quad\text{ if $v$ is in the subforest $\tau_{I}$}\\ \Delta^{\mathcal{B}}(s_{l,v,\tau_{J\cup\{I\}}}^{m})\quad\text{ otherwise.}\end{cases}$
@@
[…]
@@
7.6. Step 3. The family $\mathbfcal{A}^{\varepsilon}(C)$ is flat over ¯$\overline{\mathcal{F}}_{n}$
@@
[$\mathbb{C}[\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}]\subset\mathbb{C}[{\mathcal{W}}_{\tau}]\otimes_{\mathbb{C}[\hbar]}U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$
@@
Next, for degenerate $C$, we proceed by induction on the codimension of a stratum. According to Lemmas 7.17 and 7.18, the restrictions of $\{s_{l,v,\tau}^{m}\}$ are the generators of the same form for each of the components, embedded into $U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$ according to the subtree or subforest embedding. So the algebraic independence of the leading terms of $\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}$ follows from the argument entirely similar to that in the proof of Lemma 7.17. Namely, by [Kn], $[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}$ is a free $\Delta^{[p]}(S(\mathfrak{g})^{\mathfrak{g}})$-module and that the homomorphism $\Delta^{[p]}\cdot\operatorname{id}:S(\mathfrak{g})\otimes[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}\to S(\mathfrak{g})^{\otimes p}$ factors as $S(\mathfrak{g})\otimes[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}\to S(\mathfrak{g})\otimes_{S(\mathfrak{g})^{\mathfrak{g}}}[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}\hookrightarrow S(\mathfrak{g})^{\otimes p}$. This implies that the associated graded of the subalgebra generated by $\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}$ is the tensor product of those for smaller $n$ over their common subalgebra freely generated by some of these generators. Iterating this argument, we see that the leading terms of $\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}$ are algebraically independent for any $(C;\varepsilon)\in\mathcal{W}_{\tau}$. ∎
@@
$\mathbfcal{A}^{\varepsilon}(C)$ has constant Hilbert-Poincaré series with respect to this filtration. Thus this forms a flat family of subalgebras parametrized by ¯$\overline{\mathcal{F}}_{n}$.
@@
The last assertion follows immediately. ∎
@@
$\mathbfcal{A}^{\varepsilon}_{\chi}(C)$ has constant Hilbert-Poincaré series with respect to this filtration. Thus this forms a flat family of subalgebras parametrized by ¯$\overline{\mathcal{F}}_{n}$.
@@
Consider the standard Hermitian form on $V(\lambda)$ uniquely determined (up to a constant factor) by the property $f_{i}^{+}=e_{i},\ h_{i}^{+}=h_{i}$, where $+$ denotes the Hermitian conjugation. Let $V({\underline{\lambda}})=V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n})$ be the tensor product of irreducible finite-dimensional representations of $\mathfrak{g}$ with highest weights $\lambda_{1},\ldots,\lambda_{n}$. We consider $V({\underline{\lambda}})$ with the standard Hermitian form which is the product of standard Hermitian forms on each component. Let $V({\underline{\lambda}})^{\mathfrak{n}_{+}}\subset V({\underline{\lambda}})$ be the subspace of singular vectors.

### v2

@@ 5.5. Compactification of the family of trigonometric Gaudin subalgebras
$(U\mathfrak{g}^{\otimes m+1})^{\mathfrak{g}}$$(U\mathfrak{g}^{\otimes n+1})^{\mathfrak{g}}$$U\mathfrak{g}^{\otimes m}$$U\mathfrak{g}^{\otimes n}$$\scriptstyle{\Delta^{\mathcal{B}}}$$\scriptstyle{\psi_{\theta}}$$\scriptstyle{\psi_{\theta}}$$\scriptstyle{\Delta^{\mathcal{B}^{\prime}}}$
Figure 4. This is a curve $C\in\overline{M}_{6+2}$. We have $n=6$, $m=4$, and $\mathcal{B}=(\{0\},\{1,2\},\{3\},\{4,5,6\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}^{trig}_{\theta}(C)=\Delta^{\mathcal{B}^{\prime}}(\mathcal{A}^{trig}_{\theta}(w_{1},w_{2},w_{3}))\mathcal{A}(z_{1},z_{2})\mathcal{A}(z_{4},z_{6},z_{6})$.
@@
Indeed, the algebra on the left is the product of the two on the right since it is so for homogeneous Gaudin algebras. Next, according to [Kn] (see [R3, Lemma 3.10] for more details) the product of $\Delta^{\mathcal{B}^{\prime}}(U\mathfrak{g}^{\otimes m})$ and the subalgebra of invariants $(U\mathfrak{g}^{\otimes n})^{\Delta^{\mathcal{B}^{\prime}}\mathfrak{g}}$ inside $U\mathfrak{g}^{\otimes n}$ is the tensor product over $Z^{\prime}$ and both are free $Z^{\prime}$-modules. Since $\Delta^{\mathcal{B}^{\prime}}(\mathcal{A}^{trig}_{\theta}(w_{1},\dots,w_{m}))\subset\Delta^{\mathcal{B}^{\prime}}(U\mathfrak{g}^{\otimes m})$, $\bigotimes_{k=1}^{m}\mathcal{A}(C_{k})\subset(U\mathfrak{g}^{\otimes n})^{\Delta^{\mathcal{B}^{\prime}}\mathfrak{g}}$ and both contain $Z^{\prime}$, the first assertion follows. Finally, since both are free $Z^{\prime}$-modules, the statement about Hilbert-Poincaré series follows by multiplying the Hilbert-Poincaré series of the tensor factors and dividing by the Hilbert-Poincaré series of $Z^{\prime}$. ∎
@@
Figure 5. This is a curve $C\in\overline{M}_{6+2}$ where $z_{0}$ does not lie on the same component as $z_{6+1}$. We have $n=6$, $m=4$, and $\mathcal{B}=(\{0,1,2,3\},\{4\},\{5\},\{6\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}^{trig}_{\theta}(C)=\psi_{\theta}(j_{\mathcal{B}}^{6+1}\mathcal{A}(0,z_{1},z_{2},z_{3})))\psi_{\theta}(\Delta^{\mathcal{B}}(\mathcal{A}(0,z_{4},z_{5},z_{6}))$.
@@
We use the same idea as in the proof of Theorem 5.5 to show that the above coefficients of (15) and (16) are algebraically independent, and this would be sufficient. Namely, we consider the filtration on $S(\mathfrak{g})^{\otimes n}$ where the $n$-th component has degree $1$ and all the others have degree 0, take the leading terms with respect to this filtration, and show that the differentials of these leading terms at the point $y=(h,\ldots,h,\underbrace{f,\ldots,f}_{m-1},f+h)$ are linearly independent.
@@ 6.2. Compactification of the family of inhomogeneous Gaudin subalgebras
We will now study the compactification of the family of inhomogeneous Gaudin subalgebras with fixed $\chi\in\mathfrak{g}^{reg}$. This family will be parametrized by the space $\overline{F}_{n}$ of cactus flower curves from section 3.2. Recall that $\overline{F}_{n}$ is a compatification of $F_{n}=(\mathbb{C}^{n}\setminus\Delta)/\mathbb{C}$. For $(z_{1},\dots,z_{n})\in F_{n}$, we have the inhomogeneous Gaudin subalgebra $\mathcal{A}_{\chi}(z_{1},\dots,z_{n})$.
@@
Figure 6. This is a curve $C\in\overline{F}_{8}$. We have $n=8$, $m=3$, and ${\mathcal{S}}=(\{1,2,3,4\},\{5\},\{6,7,8\})$. The curve $C_{1}\in F_{4}$ has $\mathcal{B}=(\{1\},\{2,3,4\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}_{\chi}(C)=\Delta^{\mathcal{B}}(\mathcal{A}_{\chi}(w_{1},w_{2}))\mathcal{A}(z_{2},z_{3},z_{4})\otimes\mathcal{A}_{\chi}\otimes\mathcal{A}_{\chi}(z_{6},z_{7},z_{8})$.
@@ 6.3. Conjectural closure for arbitrary $\chi\in\mathfrak{h}$
$F_{n,\mathfrak{g}}:=((\mathbb{C}^{n}\setminus\Delta)/\mathbb{C})\times\mathfrak{h}^{reg}/\mathbb{C}^{\times}$
@@
For $n=1$ this Conjecture is [HKRW, Theorem 10.8]. Also, the particular case of this Conjecture with $n=2$ and $\mathfrak{g}_{1}$ being a root $\operatorname{\mathfrak{sl}}_{2}$ is discussed in [HKRW, Proposition 10.17]. Moreover, we expect that for $\mathfrak{g}=\operatorname{\mathfrak{sl}}_{m}$ all such spaces $\overline{F}_{n,\mathfrak{g}}$ are given by the construction of Bottman and Oblomkov [BO] of a moduli space of marked vertical lines in $\mathbb{C}^{2}$ with n1n-1 vertical lines without marked points and another vertical line with $m$ marked points [^5] We thank Alexei Oblomkov for pointing out this conjectural relation..
@@ 7.1. Degeneration of trigonometric to inhomogeneous
Recall ${\mathcal{F}}_{n}:=({\mathbb{G}}^{n}\setminus\Delta)/{\mathbb{G}}$, as defined in section 3.3. Our goal in this section is to define a family of subalgebras of $U\mathfrak{g}^{\otimes n}$ parametrized by $\mathcal{F}_{n}$ and to prove the following result.
@@
    After identifying $U_{0}\mathfrak{g}=S(\mathfrak{g})$ and $U_{\varepsilon}\mathfrak{g}=U\mathfrak{g}$, for $\varepsilon\neq 0$, this map equals $\operatorname{ev}_{\infty,z_{1},\dots,z_{n}}$ if $\varepsilon=0$ and $\operatorname{ev}_{\varepsilon^{-1},z_{1},\dots,z_{n}}$ otherwise.
@@
$(U_{\varepsilon}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n})^{\mathfrak{g}}$$U_{\varepsilon}\mathfrak{h}\otimes U\mathfrak{g}^{\otimes n}$$\mathbb{C}[\hbar]/(\hbar-\varepsilon)\otimes U\mathfrak{g}^{\otimes n}$$(U\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n})^{\mathfrak{g}}$$U\mathfrak{h}\otimes U\mathfrak{g}^{\otimes n}$$U\mathfrak{g}^{\otimes n}$$\scriptstyle{\operatorname{Rees}\psi}$$\scriptstyle{\chi}$$\scriptstyle{\psi}$$\scriptstyle{\varepsilon^{-1}\chi}$
@@ 7.2. Quadratic parts of inhomogeneous/trigonometric Gaudin subalgebras
[…]
@@
$\overline{{\mathcal{F}}}_{n}\rightarrow{\mathrm{Gr}}_{\mathbb{A}^{1}}(2n+1,U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n})\quad\overline{{\mathcal{F}}}_{n}\rightarrow{\mathrm{Gr}}_{\mathbb{A}^{1}}(2n,\mathbb{C}[\hbar]\otimes U\mathfrak{g}^{\otimes n})$
@@
In section 7.1, we constructed $\mathbfcal A$, a family of subalgebras of $U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$ parametrized by ${\mathcal{F}}_{n}$, and $\mathbfcal A_{\chi}$, a family of subalgebras of $\mathbb{C}[\hbar]\otimes U\mathfrak{g}^{\otimes n}$ also parametrized by ${\mathcal{F}}_{n}$ (in both cases regarded over the base ring $\mathbb{C}[\hbar]$). Recall also that in [IKLPR], we constructed a (relative) compactification $\overline{{\mathcal{F}}}_{n}$, see section 3.3.
@@
These extend to flat families of subalgebras faithfully parametrized by $\overline{{\mathcal{F}}}_{n}$. Thus $\overline{{\mathcal{F}}}_{n}$ is the compactified parameter space for these inhomogeneous/trigonometric subalgebras.
As in the proof of Theorem 5.18, it suffices to show that the family extends to a flat family parametrized by $\overline{{\mathcal{F}}}_{n}$. Subsequently, the faithfulness of the parametrization follows from Corollary 7.10.
@@
    Each ${\mathcal{W}}_{\tau}$ is a union of strata of $\overline{{\mathcal{F}}}_{n}$. These strata are constructed in an “operadic manner” from smaller $\overline{{\mathcal{F}}}_{r}$ and $\overline{M}_{k+1}$. We prove that the restrictions of our generators to these strata are compatible with the generators of the Gaudin subalgebras over these strata.
@@
On the other hand, the sum of residues of $f(w)$ at all $w_{i}$ and $\infty$ is 0, so we have
@@
Let $\tau$ be a binary rooted forest. Let $v$ be an internal vertex of $\tau$. Let $p$ be the rightmost leaf of the left branch growing from $v$ and let $q$ be the leftmost leaf on the right branch growing from $v$. We make an affine change of the variable $u$ taking $z_{p}$ to 0 and $z_{q}$ to $1$; the new variable is
@@
Now, suppose that $v$ is a root of $\tau$. Let $p$ be the rightmost leaf on the tree growing from $v$. We make an affine change of the variable $u$ taking $z_{p}$ to 0 and preserving $\hbar^{-1}$; the new variable is
@@
In [IKLPR, §6.3], we considered a stratification of $\overline{{\mathcal{F}}}_{n}$ indexed by pairs of set partitions; on each stratum certain of the $\nu_{ij}$ coordinates are equal to 0 and others are equal to $\infty$ (as $\delta_{ij}=\nu_{ij}^{-1}$, this is equivalent to setting some $\delta_{ij}=0$).
Let $A$ be a codimension 1 stratum of $\overline{{\mathcal{F}}}_{n}$ such that $A\cap{\mathcal{W}}_{\tau}\neq\emptyset$. From [IKLPR, Remark 6.13], we have the following 2 possibilities:
@@
    $A=\overline{{\mathcal{F}}}_{I}\times_{\mathbb{A}^{1}}\overline{{\mathcal{F}}}_{J}$, determined by a partition $n]=I\sqcup J$ compatible the forest $\tau$, by putting $\nu_{ji}=0$ for all $i\in I,\ j\in J$;
@@
[…]
@@
Now it remains to show that the above generators stay algebraically independent. This follows from the particular case of Knop’s theorem [Kn] (see [R3, Lemma 3.10] for more details), namely $U\mathfrak{g}^{\otimes I}]^{\mathfrak{g}}$ is a free $\Delta(ZU\mathfrak{g})$-module and that the homomorphism $\Delta\cdot\operatorname{id}:U\mathfrak{g}\otimes[U\mathfrak{g}^{\otimes I}]^{\mathfrak{g}}\to U\mathfrak{g}^{\otimes I}$ factors as
@@
This implies that the product of $j_{\{0\}\cup I}^{n+1}(\mathbfcal{A}^{\varepsilon}(C_{I}))$ and $\Delta^{\mathcal{B}}(\mathbfcal{A}^{\varepsilon}(C_{J}))$ is in fact the tensor product
@@
[…]
@@
$j_{I}^{n+1}(\mathcal{A}(C_{I}))\otimes_{\Delta^{\mathcal{B}}(ZU\mathfrak{g})}\Delta^{\mathcal{B}}(\mathbfcal{A}^{\varepsilon}(C_{J\cup\{I\}}))$
@@
7.6. Step 3. The family $\mathbfcal{A}^{\varepsilon}(C)$ is flat over $\overline{\mathcal{F}}_{n}$
@@
$\mathbb{C}[\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}]\subset\mathbb{C}[{\mathcal{W}}_{\tau}]\otimes_{\mathbb{C}[\hbar]}U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$
@@
Next, for degenerate $C$, we proceed by induction on the codimension of a stratum. According to Lemmas 7.17 and 7.18, the restrictions of $\{s_{l,v,\tau}^{m}\}$ are the generators of the same form for each of the components, embedded into $U_{\hbar}\mathfrak{g}\otimes U\mathfrak{g}^{\otimes n}$ according to the subtree or subforest embedding. So the algebraic independence of the leading terms of $\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}$ follows from the argument entirely similar to that in the proof of Lemma 7.17. Namely, by [Kn], $S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}$ is a free $\Delta^{[p]}(S(\mathfrak{g})^{\mathfrak{g}})$-module and that the homomorphism $\Delta^{[p]}\cdot\operatorname{id}:S(\mathfrak{g})\otimes[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}\to S(\mathfrak{g})^{\otimes p}$ factors as $S(\mathfrak{g})\otimes[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}\to S(\mathfrak{g})\otimes_{S(\mathfrak{g})^{\mathfrak{g}}}[S(\mathfrak{g})^{\otimes p}]^{\mathfrak{g}}\hookrightarrow S(\mathfrak{g})^{\otimes p}$. This implies that the associated graded of the subalgebra generated by $\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}$ is the tensor product of those for smaller $n$ over their common subalgebra freely generated by some of these generators. Iterating this argument, we see that the leading terms of $\{s_{l,v,\tau}^{m}\}\cup\{\widetilde{\Phi}^{(0)}_{l}\}$ are algebraically independent for any $(C;\varepsilon)\in\mathcal{W}_{\tau}$. ∎
@@
$\mathbfcal{A}^{\varepsilon}(C)$ has constant Hilbert-Poincaré series with respect to this filtration. Thus this forms a flat family of subalgebras parametrized by $\overline{{\mathcal{F}}}_{n}$.
@@
The last assertion follows immediately.
@@
$\mathbfcal{A}^{\varepsilon}_{\chi}(C)$ has constant Hilbert-Poincaré series with respect to this filtration. Thus this forms a flat family of subalgebras parametrized by $\overline{{\mathcal{F}}}_{n}$.
@@
Consider the standard Hermitian form on $V(\lambda)$ uniquely determined (up to a constant factor) by the property $f_{i}^{+}=e_{i},\ h_{i}^{+}=h_{i}$, where ++ denotes the Hermitian conjugation. Let $V({\underline{\lambda}})=V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n})$ be the tensor product of irreducible finite-dimensional representations of $\mathfrak{g}$ with highest weights $\lambda_{1},\ldots,\lambda_{n}$. We consider $V({\underline{\lambda}})$ with the standard Hermitian form which is the product of standard Hermitian forms on each component. Let $V({\underline{\lambda}})^{\mathfrak{n}_{+}}\subset V({\underline{\lambda}})$ be the subspace of singular vectors.